OpenStudy (anonymous):

The length of a rectangle is 2 m less than twice the width. If the perimeter is 11 m, find the length and width.

OpenStudy (anonymous):

i don't get this

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

help me PLZZZZZ

ganeshie8 (ganeshie8):

lol stop, il help

OpenStudy (anonymous):

thank !

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

I hate word problem but my teacher make me do it

ganeshie8 (ganeshie8):

The length of a rectangle is 2 m less than twice the width. \(\large l = 2w - 2 \) --------------(1)

ganeshie8 (ganeshie8):

im ex hater of word problems too ;)

OpenStudy (anonymous):

I want the equation

ganeshie8 (ganeshie8):

If the perimeter is 11 m, \(\large 2(l+w) = 11 \) ---------(2)

ganeshie8 (ganeshie8):

you have two equations, solve them and find length and width

OpenStudy (anonymous):

2(l+w)=11 is that the equation

OpenStudy (anonymous):

i don't get it

ganeshie8 (ganeshie8):

perimeter of rectangle = 2(l+w)

OpenStudy (anonymous):

so what is the equation

OpenStudy (anonymous):

so what is the equation

OpenStudy (anonymous):